OpenStudy (anonymous):

function operations. f/g and f(x)=2x g(x)=radicalx-1

OpenStudy (anonymous):



OpenStudy (anonymous):

the -1 is NOT under the radical

OpenStudy (turingtest):

what you have so far for f/g is correct what is your question?

OpenStudy (anonymous):

well...am I done?

OpenStudy (turingtest):

sure unless you wnat to "rationalize the denominator"\[[2x/(\sqrt{x}-1)][(\sqrt{x}+1)/(\sqrt{x}+1)]=2x( \sqrt{x}+1)/(x-1)\]sometimes they want that but that's really not very simplifying in this case, so I would say you're done as you are.